### v1

# Dymnikova black hole from an infinite tower of higher-curvature corrections

Abstract
Recently, in Bueno:2024dgm , it was demonstrated that various regular black hole metrics can be derived within a theory featuring an infinite number of higher curvature corrections to General Relativity. Moreover, truncating this infinite series at the first few orders already yields a reliable approximation of the observable characteristics of such black holes Konoplya:2024hfg . Here, we further establish the existence of another regular black hole solution, particularly the $D$-dimensional extension of the Dymnikova black hole, within the equations of motion incorporating an infinite tower of higher-curvature corrections. This solution is essentially nonperturbative in the coupling parameter, rendering the action, if it exists, incapable of being approximated by a finite number of powers of the curvature.
pacs:
04.30.Nk,04.50.Kd,04.70.Bw
I Introduction
The singularity problem for the classical Schwarzschild solution marks the limits of General Relativity. While a great number of approaches to solve this problem was suggested, recently a large class of regular black holes in $D\geq 5$ spacetime dimensions have been obtained as a result of adding the infinite number of higher-curvature terms to the Einstein action Bueno:2024dgm . This approach is a part of the quasi-topological gravity Oliva:2010eb ; Myers:2010ru ; Dehghani:2011vu ; Ahmed:2017jod ; Cisterna:2017umf , so that no exotic or other badly motivated matter is introduced to produce a regular black hole. This approach has been recently studied in DiFilippo:2024mwm and quasinormal modes of regular black holes obtained within this approach have been analyzed in Konoplya:2024hfg . An important observation of Konoplya:2024hfg consisted in the justification of the lower order expansions which could serve as good approximations for observable quantities for the full regular black hole solutions.
The freedom to establish various relations between the coupling parameters allowed to produce classes of regular black-hole metrics with desired features DiFilippo:2024mwm . The natural question is then whether it is indeed so, and any regular black hole metric can be a solution to some theory which could represented as an infinite series of higher curvature corrections and, first of all, whether such a theory corresponds to some particular choice of the coupling parameters of the quasi-topological gravity?
For this purpose we will study what kind of higher curvature corrections correspond to the $D$-dimensional generalization of the Dymnikova regular black hole metric Dymnikova:1992ux . Dymanikova black hole can be interpreted as a solution of the Einstein equations (four or higher dimensional) with an energy-momentum tensor, corresponding to the exponentially decaying matter Dymnikova:1992ux ; Paul:2023pqn .
II Generalized Dymnikova metric derivation
Bueno, Cano, and Hennigar Bueno:2024dgm considered the action of the Einsteinian theory with general higher curvature corrections
$I_{\rm QT}=\frac{1}{16\pi G}\int\mathrm{d}^{D}x\sqrt{|g|}\left[R+\sum_{n=2}^{n_{\rm max}}\alpha_{n}\mathcal{Z}_{n}\right]\,,$ (1)
where $\alpha_{n}$ are arbitrary coupling constants with dimensions of length2(n-1) and $\mathcal{Z}_{n}$ are the Quasi-topological densities Bueno:2019ycr . Regular black-hole solutions correspond to theories with the infinite number of terms in the above action.
The line element for the spherically symmetric $D$-dimensional black hole is
$ds^{2}$ $=$ $-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega^{2}_{D-2}\,,$
$f(r)\equiv 1-r^{2}\psi(r)\,.$
The equations of motion imply that
$\dfrac{d}{dr}\left(r^{D-1}h(\psi(r))\right)=0,$ (3)
where the function $h(\psi)$ is given by the series,
$h(\psi)\equiv\psi+\sum_{n=2}^{n_{\rm max}}\alpha_{n}\psi^{n}\,.$ (4)
The solution to Eq. 3 is
$h(\psi(r))=\dfrac{\mu}{r^{D-1}}\,,$ (5)
where $\mu$ is the positive constant proportional to the ADM mass.
In Bueno:2024dgm a limit of indefinite tower of higher curvature corrections in (1) is introduced via an explicit function $h(\psi)$, which is monotonously growing for $0<\psi<\psi_{0}$ and diverges at $\psi=\psi_{0}$. In this way, the solution to the equation (5) gives a black-hole configuration, in which, by construction, the singularity at $r=0$ is replaced by the de Sitter core,
$f(r)=1+{\cal O}\left(r^{2}\right)\approx 1-r^{2}\psi_{0}.$ (6)
Since the function $h(\psi)$ is initially defined through the series (4), it is implicitly assumed that $h(\psi)$ is regular in $\psi=0$. We note that regularity in this point is not a necessary condition. The reason for that is that $\psi=0$ corresponds to the asymptotic infinity of the spacetime, which is also not a regular point. Moreover, in the present paper we propose a simple example of such a function $h(\psi)$, which is not regular in $\psi=0$, while leading to a meaningful black-hole configuration free from singularities.
A simple function of $\psi\geq 0$, which is not regular at $\psi=0$, is $e^{-1/\alpha\psi}$, where $\alpha$ is a positive constant. This is a monotonous analytic function for $\psi>0$, which cannot be expanded in Taylor series near the origin. Keeping in mind this feature, we are in a position to construct the function $h(\psi)$, which also diverges at $\psi=\psi_{0}\equiv\alpha^{-1}>0$. We define the function as follows:
$h(\psi)\equiv\dfrac{\psi}{1+\alpha\psi W_{0}\left(-\dfrac{e^{-1/\alpha\psi}}{\alpha\psi}\right)},\quad\alpha>0,$ (7)
where $W_{0}(z)$ is the principal Lambert function, which satisfies the equation
$W_{0}(z)e^{W_{0}(z)}=z,$
and
$W_{0}(0)=0,\quad W_{0}(-e^{-1})=-1.$
For the function $h(\psi)$ given by (7) the solution to the Eq. (5) can be found in an explicit form,
$\psi(r)=\dfrac{\mu}{r^{D-1}}\left(1-e^{-\dfrac{r^{D-1}}{\alpha\mu}}\right),$ (8)
and the corresponding metric function is
$f(r)=1-\dfrac{\mu}{r^{D-3}}\left(1-e^{-\dfrac{r^{D-1}}{\alpha\mu}}\right).$ (9)
The metric (9) is analytic for $r\geq 0$. It coincides with the one obtained in higher dimensional Einstein theory in the presence of matter Paul:2023pqn and, when $D=4$, it is reduced to the well-known Dymnikova black hole Dymnikova:1992ux ; Platania:2019kyx .
It is important to stress out that $h(\psi)$ cannot be represented as series (4), and consequently, the action cannot be represented as (1). Although the limit of $\alpha\to 0$ clearly exists and is the Schwarzschild solution to General Relativity, the theory is not perturbative in terms of the quasi-topological corrections. Consequently, one is unable to approximate such a theory with a finite number of the higher-order corrections.
III Discussion
Although we are unable to represent the action of the theory in the form (4), it does not mean that such a theory does not exist. In order to illustrate this point we can propose an ad hoc action with a similar feature
$I=\frac{1}{16\pi G}\int\mathrm{d}^{D}x\sqrt{|g|}\left(R+\dfrac{1}{\alpha}e^{-\dfrac{1}{\alpha^{2}R^{2}}}\right)\,.$ (10)
This theory is nonperturbative in terms of the coupling constant $\alpha$ and cannot be approximated by finite number of the higher-order curvature corrections. At the same time, it has well-defined equations of motion and approaches General Relativity in the limit of $\alpha\to 0$. Thus, we cannot exclude that there is a well-defined theory of gravity, in which a $D$-dimensional Dymnikova black hole is a vacuum solution. We conclude that this theory must be nonperturbative in terms of the higher-order curvature corrections.
The metric function (9) cannot be expanded in the series of $\alpha$ either and, unlike the cases considered in Konoplya:2024hfg , one cannot calculate the quasinormal modes using the Frobenius method applied to such an approximate metric. In order to be able to apply the Frobenius method to the Dymnikova black hole it was proposed in Konoplya:2023aph to use the parametrization Rezzolla:2014mua , which was generalized for $D>4$ in Konoplya:2020kqb .
Apparently, the analysis of the quasinormal ringing in this case can be done using the methods, which do not involve the analysis of the asymptotic behaviour of the metric functions. Thus, the WKB formula Konoplya:2019hlu has recently been used for calculations of the quasinormal modes of the scalar field in Macedo:2024dqb . Another potentially useful approach would be employing the spectral methods for numerically solving the eigenvalue problem Jansen:2017oag ; Fortuna:2020obg ; Konoplya:2022zav .